### v1

@@ 5 Research Question Results
    Summary of RQ1: Fine-tuned by a dataset of file-level secure code practices, LLMs show marginal improvements in generating secure code. Through a qualitative analysis, we find that the fine-tuned LLMs might learn both vulnerability fix patterns and new semantic patterns that might be hard to generalize, hence resulting in more invalid code but also more secure code.
    RQ2: How the granularity of fine-tuning dataset may impact fine-tuning LLMs for secure code generation?
    Motivation. RQ1 investigates only one version of the fine-tuning dataset, i.e., source code files that contain vulnerability fixes. RQ1 findings motivate us to explore diverse versions of the same fine-tuning datasets. For instance, fine-tuned LLMs may learn incomplete semantics (“noise”) from the fine-tuning dataset, which results in a larger number of invalid code compared to the pre-trained LLMs. Using finer-grained datasets for fine-tuning may reduce the confusion introduced by fine-tuning examples and allow LLMs to focus on learning vulnerability fix patterns.
    Approach. We crafted three additional fine-tuning datasets based on granularity from the same set of 4,900 vulnerability-fixing commits. The one used in RQ1 is file-level. In this RQ, we created function-level, block-level, and line-level fine-tuning datasets. The file-level dataset consists of all the files that have modifications in the vulnerability-fixing commits. For the function-level dataset, we first located the modified lines in each commit and then extracted only the functions containing these modified lines. Similarly, we collected the block-level dataset by locating the nearest blocks containing the modified lines. Finally, the line-level dataset consists of all the modified lines in the vulnerability-fixing commits. Note that the four granularities of the fine-tuning sets are all from the same 4,900 commits. We extract all the blocks, functions, or lines relevant to the changed lines for each commit. This RQ aims to explore the impact of different code contexts on fine-tuning performance. Table 4 summarizes the size of the four datasets of different granularity. After constructing the four datasets, we apply these datasets to fine-tune CodeLlama, and CodeGen2 using the two parameter-efficient fine-tuning techniques.
    Granularity Total Number of Items Used for Fine-tuning
    file-level 14,622 files
    function-level 32,244 functions
    block-level 79,464 blocks
    line-level 1,240,407 lines
    Results. Table 3 shows the vulnerability evaluation results of the four fine-tuned models in different granularity. In general, we notice that the LLMs fine-tuned by finer-grained (block, function, and line) datasets usually perform better than the file-level (in 7 out of 8 testing settings).
    The block-level fine-tuned models perform better on the generation of secure code snippets compared to other models fine-tuned from other granularities (i.e., it achieves the best performance in 4 out of 8 testing settings). LoRA achieved the highest secure ratio in C language with 62.4% for CodeGen2-7B with a 6.4% increase compared with the pre-trained model. Another model, CodeLlama-7B, has a 60.7% secure ratio. In C++, models fine-tuned from function-level data outperform the block-level ones, with a 64.2% secure ratio for CodeGen2-7B and 65.1% for CodeLlama-7B. However, full-weight fine-tuning results in a decrease in valid ratio, averaging a 10% drop in C and a 24% drop in C++. Similarly, parameter-efficient fine-tuning techniques encountered a drop in terms of valid code generation in C++.
    (a) The comparison of the pre-trained models and the fine-tuned models with the highest secure ratio in C language for each CWE.
    (b) The comparison of the pre-trained models and the fine-tuned models with the highest secure ratio in C++ language for each CWE.
    Figure 6: Secure ratio of fine-tuned LLMs for the 11 CWEs.
    Summary of RQ2: Finer-grained vulnerability-fixing code often results in fine-tuned LLMs with better performance, as evidenced by a higher secure ratio. This suggests that finer-grained datasets may compel LLMs to learn vulnerability fixing patterns more effectively by excluding irrelevant code.
    RQ3: How the size of fine-tuning dataset may impact fine-tuning LLMs for secure code generation?
    Motivation. The size of our fine-tuning dataset may affect the fine-tuning process. It is unknown how much impact the fine-tuning dataset of different sizes has. Thus, we try to uncover the influence of different dataset sizes in terms of secure code generation in this RQ.
@@
    Valid code/Total Secure code/Valid code Valid code/Total Secure code/Valid code
@@
    100 commits 80/90 (88.9%) 50/80 (62.5%) 64/90 (71.1%) 52/64 (81.2%)
    200 commits 76/90 (84.4%) 55/76 (72.4%) 59/90 (65.6%) 44/59 (74.6%)
    300 commits 79/90 (87.8%) 60/79 (75.9%) 66/90 (73.3%) 52/66 (78.8%)
    400 commits 73/90 (81.1%) 59/73 (80.8%) 66/90 (73.3%) 49/66 (74.2%)
    500 commits 72/90 (80.0%) 60/72 (83.3%) 62/90 (68.9%) 46/62 (74.2%)
    All commits 77/90 (85.6%) 61/77 (79.2%) 72/90 (80.0%) 45/72 (62.5%)
@@
    100 commits 81/90 (90.0%) 53/81 (65.4%) 71/90 (78.9%) 44/71 (62.0%)
    200 commits 79/90 (87.8%) 50/79 (63.3%) 65/90 (72.2%) 38/65 (58.5%)
    300 commits 77/90 (85.6%) 53/77 (68.8%) 79/90 (87.8%) 53/79 (67.1%)
    400 commits 77/90 (85.6%) 48/77 (62.3%) 74/90 (82.2%) 46/74 (62.2%)
    500 commits 78/90 (86.7%) 51/78 (65.4%) 76/90 (84.4%) 57/76 (75.0%)
    All commits 77/90 (85.6%) 54/77 (70.1%) 72/90 (80.0%) 47/72 (65.3%)
@@
    100 commits 71/90 (78.9%) 55/71 (77.5%) 56/90 (62.2%) 42/56 (75.0%)
    200 commits 74/90 (82.2%) 62/74 (83.8%) 70/90 (77.8%) 51/70 (72.9%)
    300 commits 71/90 (78.9%) 54/71 (76.1%) 64/90 (71.1%) 52/64 (81.2%)
    400 commits 69/90 (76.7%) 53/69 (76.8%) 70/90 (77.8%) 61/70 (87.1%)
    500 commits 70/90 (77.8%) 50/70 (71.4%) 56/90 (62.2%) 37/56 (66.1%)
    All commits 72/90 (80.0%) 53/72 (73.6%) 61/90 (67.8%) 51/61 (83.6%)
@@
    100 commits 79/90 (87.8%) 48/79 (60.8%) 73/90 (81.1%) 56/73 (76.7%)
    200 commits 80/90 (88.9%) 59/80 (73.8%) 70/90 (77.8%) 50/70 (71.4%)
    300 commits 72/90 (80.0%) 51/72 (70.8%) 76/90 (84.4%) 57/76 (75.0%)
    400 commits 78/90 (86.7%) 60/78 (76.9%) 75/90 (83.3%) 55/75 (73.3%)
    500 commits 78/90 (86.7%) 59/78 (75.6%) 77/90 (85.6%) 57/77 (74.0%)
    All commits 76/90 (84.4%) 51/76 (67.1%) 81/90 (90.0%) 55/81 (67.9%)
    Approach. As our fine-tuning dataset contains fixes of multiple CWEs, it is difficult to evaluate the different sizes in different CWEs in our fine-tuning dataset. Thus, we pick up one CWE and collect all fixes commits from our fine-tuning dataset. CWE-119 has the most commits (507 commits) in our dataset compared with other CWE. After collecting all CWE-119 commits, we take a filtering process on them to remove some fix-irrelevant (e.g., refactoring and comment modifications) files. There are three commits that have too many file changes and are hard to find the fix files, so we remove these commits. Overall, we craft a CWE-119 dataset with 507 commits, including 997 files.
    To uncover the impact of different sizes of datasets in fine-tuning process. We construct different sizes of the CWE-119 dataset. Particularly, we first randomly sample files of 100 commits from the whole CWE-119 dataset as dataset1, Then, we keep the files of dataset combined with another 100 sampled commits from the rest of the commits as dataset2. We repeat this process until all commits are sampled. Finally, six datasets were created (Table 5).
    In the end, we fine-tuned models using these datasets of different sizes individually with 4 epochs.
    Results. Table 6 presents the results in different sizes of fine-tuning datasets in CodeGen2 and CodeLlama. CWE-119 has three scenarios. For each scenario, we asked each model to generate 30 code snippets, resulting in 90 generated snippets per setting. It was observed that most models generate more valid cases with the increase in dataset size (in 5 out of 8 settings).
    In terms of C language, CodeGen2 using IA3 in 100 commits achieved the highest valid ratio at 90%. Similarly, with a larger number of commits in the training set, CodeGen2 achieved a higher secure ratio (i.e., 83.3% for 500 commits and 79.2% for all commits in LoRa and 70.1% for all commits with IA3). CodeGen2 fine-tuned on all commits is higher than the one fine-tuned on 100 commits in the secure ratio for both LoRa and IA3. CodeLlama using IA3 in 200 commits achieved the highest valid ratio (88.9%). The model trained with 200 commits in LoRa is the highest secure ratio at 83.8%.
    For C++ language, CodeGen2, using the LoRa fine-tuning method, the valid code ratios range from 65.6% to 80.0%, with the highest valid code ratio being 80.0% at all commits and the lowest being 65.6% at 200 commits. The secure ratios for valid code vary between 62.5% and 81.2%. The highest valid ratio is in CodeGen2 using IA3 at 300 comments, and the ratio is 87.8%. The highest secure ratio is in CodeGen2 using LoRa at 100 commits, which is 81.2%. CodeLlama using IA3 at all commits achieved the highest valid ratio at 90.0%. However, it also got the lowest secure ratio with 67.9% compared to other sizes of training sets. The CodeLlama using LoRa at 400 commits achieved the highest secure ratio at 87.1%, and its valid ratio is relatively higher than other training sets.
    Summary of RQ3: The size of the fine-tuning dataset does have a non-trivial impact on the improvement of generating secure code. While a larger fine-tuning dataset may let LLMs generate less vulnerable code, it may produce more invalid code.
@@
    Approach. Since we focus on C/C++ in this study, we selected tasks of C++ language from HumanEval-X. We call it as HumanEval_CPP. For each task, we generated 10 samples for each pre-trained and fine-tuned version of CodeGen2 and CodeLlama. Then, we calculate PASS@1, meaning that for one prompt, whether or not an LLM can generate one correct code given one attempt.
    Results. Table 7 presents PASS@1 results of all our fine-tuned LLMs on HumanEval_CPP dataset. CodeGen2 suffers a slight performance degradation, especially when fine-tuned with IA3. CodeLlama shows improvements when fine-tuned by IA3, and comparable results (no performance degradation) when fine-tuned by LoRA. When using file-level in CodeGen2, it achieved 9.09% and 9.33% in LoRA and IA3, respectively, which are close to the performance of the pre-trained model. Additionally, we conducted a Fisher exact test to compare the code generation results between one original LLM and one fine-tuned LLM for all the 164 tasks in the HumanEval dataset. Our results show that for all the LLMs in our study, no significant difference exists between the original and the fine-tuned LLMs. For most of the 164 tasks, the P-value is greater than 0.05, which means that it does not affect the functional correctness of the code generation between the pre-trained model and the fine-tuned model. To give an example, comparing CodeLLama and the fine-tuned CodeLLama using LORA and file-level fine-tuning dataset, for only two of 164 tasks in HumanEval, we observed a statically significant difference.
@@
    CodeGen2-7B Line-level 8.05% CodeLlama-7B Line-level 10.73%
@@
    File-level 9.33% File-level 12.87%

### v2

@@ 5 Research Question Results
    - (4) How the size of fine-tuning dataset may impact fine-tuning LLMs for secure code generation? The size of our fine-tuning dataset may affect the fine-tuning process. We try to uncover the influence of different dataset sizes in terms of secure code generation. As our fine-tuning dataset contains fixes of multiple CWEs, it is difficult to evaluate the different sizes in different CWEs in our fine-tuning dataset. Thus, we pick up one CWE and collect all fixes commits from our fine-tuning dataset. CWE-119 has the most commits (507 commits) in our dataset compared with other CWE. Overall, we craft a CWE-119 dataset with 507 commits, including 997 files. To uncover the impact of different sizes of datasets in the fine-tuning process. We construct different sizes of the CWE-119 dataset. Particularly, we first randomly sample files of 100 commits from the whole CWE-119 dataset as dataset1, Then, we keep the files of dataset combined with another 100 sampled commits from the rest of the commits as dataset2. We repeat this process until all commits are sampled. Finally, six datasets were created (Table 4). Table 5 presents the results in different sizes of fine-tuning datasets in four models. CWE-119 has three scenarios. For each scenario, we asked each model to generate 30 code snippets, resulting in 90 generated snippets per setting. It was observed that most models generate more valid cases with the increase in dataset size (in 10 out of 16 settings compared between the 100 commits and the all commits). In terms of C, CodeGen2 using IA3 in 100 commits achieved the highest valid ratio at 90%. Similarly, with a larger number of commits in the training set, CodeGen2 achieved a higher secure ratio (i.e., 83.3% for 500 commits and 79.2% for all commits in LoRA and 70.1% for all commits with IA3). CodeGen2 fine-tuned on all commits is higher than the one fine-tuned on 100 commits in the secure ratio for both LoRA and IA3. CodeLlama using IA3 in 200 commits achieved the highest valid ratio (88.9%). The model trained with 200 commits in LoRA is the highest secure ratio at 83.8%. DeepSeek-Coder achieved the highest secure ratios on 69.3% using LoRA with the dataset of 200 commits and 67.8% using IA3 with the dataset of 300 commits. Using the dataset of all commits, Magicoder fine-tuned by LoRA achieved 75.3%, the highest secure ratio among other datasets. For C++ language, CodeGen2, using the LoRA fine-tuning method, the valid code ratios range from 65.6% to 80.0%, with the highest valid code ratio being 80.0% at all commits and the lowest being 65.6% at 200 commits. The secure ratios for valid code vary between 62.5% and 81.2%. The highest valid ratio is in CodeGen2 using IA3 at 300 commits, and the ratio is 87.8%. The highest secure ratio is in CodeGen2 using LoRA at 100 commits, which is 81.2%. CodeLlama using IA3 at all commits achieved the highest valid ratio at 90.0%. However, it also got the lowest secure ratio with 67.9% compared to other sizes of training sets. The CodeLlama using LoRA at 400 commits achieved the highest secure ratio at 87.1%, and its valid ratio is relatively higher than other training sets. The highest secure ratios in LoRA and IA3 of DeepSeek-Coder are 75.9% and 74.1% respectively. In terms of Magicoder, the dataset of all commits achieved the highest secure ratio with 87.8% using LoRA.
@@
    Valid code / Secure code / Valid code / Secure code /
    Total Valid code Total Valid code
@@
    100 commits 80/90 (88.9%) 50/80 (62.5%) 73/90 (81.1%) 55/73 (75.3%)
    200 commits 76/90 (84.4%) 55/76 (72.4%) 82/90 (91.1%) 61/82 (74.4%)
    300 commits 79/90 (87.8%) 60/79 (75.9%) 84/90 (93.3%) 63/84 (75.0%)
    400 commits 73/90 (81.1%) 59/73 (80.8%) 77/90 (85.6%) 60/77 (77.9%)
    500 commits 72/90 (80.0%) 60/72 (83.3%) 76/90 (84.4%) 55/76 (72.4%)
    All commits 77/90 (85.6%) 61/77 (79.2%) 80/90 (88.9%) 50/80 (62.5%)
@@
    100 commits 81/90 (90.0%) 53/81 (65.4%) 82/90 (91.1%) 50/82 (61.0%)
    200 commits 79/90 (87.8%) 50/79 (63.3%) 78/90 (86.7%) 47/78 (60.3%)
    300 commits 77/90 (85.6%) 53/77 (68.8%) 81/90 (90.0%) 54/81 (66.7%)
    400 commits 77/90 (85.6%) 48/77 (62.3%) 83/90 (92.2%) 53/83 (63.9%)
    500 commits 78/90 (86.7%) 51/78 (65.4%) 88/90 (97.8%) 63/88 (71.6%)
    All commits 77/90 (85.6%) 54/77 (70.1%) 81/90 (90.0%) 54/81 (66.7%)
@@
    100 commits 71/90 (78.9%) 55/71 (77.5%) 73/90 (81.1%) 52/73 (71.2%)
    200 commits 74/90 (82.2%) 62/74 (83.8%) 81/90 (90.0%) 57/81 (70.4%)
    300 commits 71/90 (78.9%) 54/71 (76.1%) 75/90 (83.3%) 60/75 (80.0%)
    400 commits 69/90 (76.7%) 53/69 (76.8%) 83/90 (92.2%) 67/83 (80.7%)
    500 commits 70/90 (77.8%) 50/70 (71.4%) 73/90 (81.1%) 47/73 (64.4%)
    All commits 72/90 (80.0%) 53/72 (73.6%) 73/90 (81.1%) 60/73 (82.2%)
@@
    100 commits 79/90 (87.8%) 48/79 (60.8%) 84/90 (93.3%) 62/84 (73.8%)
    200 commits 80/90 (88.9%) 59/80 (73.8%) 83/90 (92.2%) 60/83 (72.3%)
    300 commits 72/90 (80.0%) 51/72 (70.8%) 85/90 (94.4%) 63/85 (74.1%)
    400 commits 78/90 (86.7%) 60/78 (76.9%) 86/90 (95.6%) 62/86 (72.1%)
    500 commits 78/90 (86.7%) 59/78 (75.6%) 85/90 (94.4%) 64/85 (75.3%)
    All commits 76/90 (84.4%) 51/76 (67.1%) 86/90 (95.6%) 57/86 (66.3%)
    LoRa
    100 commits 86/90 (95.6%) 58/86 (67.4%) 87/90 (96.7%) 63/87 (72.4%)
    200 commits 90/90 (100.0%) 59/90 (65.6%) 87/90 (96.7%) 64/87 (73.6%)
    300 commits 88/90 (97.8%) 61/88 (69.3%) 88/90 (97.8%) 58/88 (65.9%)
    400 commits 87/90 (96.7%) 60/87 (69.0%) 89/90 (98.9%) 65/89 (73.0%)
    500 commits 89/90 (98.9%) 54/89 (60.7%) 87/90 (96.7%) 66/87 (75.9%)
    All commits 90/90 (100.0%) 59/90 (65.6%) 88/90 (97.8%) 62/88 (70.5%)
    DeepSeek-Coder-7B IA3
    100 commits 89/90 (98.9%) 60/89 (67.4%) 89/90 (98.9%) 61/89 (68.5%)
    200 commits 90/90 (100.0%) 61/90 (67.8%) 87/90 (96.7%) 56/87 (64.4%)
    300 commits 89/90 (98.9%) 52/89 (58.4%) 85/90 (94.4%) 63/85 (74.1%)
    400 commits 89/90 (98.9%) 53/89 (59.6%) 88/90 (97.8%) 60/88 (68.2%)
    500 commits 89/90 (98.9%) 59/89 (66.3%) 89/90 (98.9%) 60/89 (67.4%)
    All commits 89/90 (98.9%) 60/89 (67.4%) 86/90 (95.6%) 53/86 (61.6%)
    LoRa
    100 commits 87/90 (96.7%) 57/87 (65.5%) 88/90 (97.8%) 69/88 (78.4%)
    200 commits 88/90 (97.8%) 61/88 (69.3%) 88/90 (97.8%) 70/88 (79.5%)
    300 commits 87/90 (96.7%) 55/87 (63.2%) 86/90 (95.6%) 72/86 (83.7%)
    400 commits 87/90 (96.7%) 61/87 (70.1%) 90/90 (100.0%) 76/90 (84.4%)
    500 commits 87/90 (96.7%) 56/87 (64.4%) 88/90 (97.8%) 68/88 (77.3%)
    All commits 89/90 (98.9%) 67/89 (75.3%) 90/90 (100.0%) 79/90 (87.8%)
    Magicoder-6.7B IA3
    100 commits 90/90 (100.0%) 57/90 (63.3%) 89/90 (98.9%) 55/89 (61.8%)
    200 commits 89/90 (98.9%) 55/89 (61.8%) 90/90 (100.0%) 56/90 (62.2%)
    300 commits 90/90 (100.0%) 48/90 (53.3%) 89/90 (98.9%) 58/89 (65.2%)
    400 commits 90/90 (100.0%) 55/90 (61.1%) 87/90 (96.7%) 56/87 (64.4%)
    500 commits 90/90 (100.0%) 50/90 (55.6%) 90/90 (100.0%) 55/90 (61.1%)
    All commits 90/90 (100.0%) 51/90 (56.7%) 90/90 (100.0%) 54/90 (60.0%)
    Summary of RQ1: Fine-tuned by a dataset of block-level secure code practices, LLMs show marginal improvements in generating secure code. Full fine-tuning can improve the secure ratio while it generated more invalid code snippets. Through a qualitative analysis, we find that the fine-tuned LLMs might learn both vulnerability fix patterns and new semantic patterns that might be hard to generalize, hence resulting in more invalid code but also more secure code. Furthermore, we observe that increasing the amount of training data generally leads to a higher secure ratio.
    RQ2: How does fine-tuning compare with prompt-based methods?
    Motivation. RQ1 explored the different PEFT techniques in fine-tuning models. It is unclear the secure performance of LLMs using prompt-based approaches. This RQ evaluates the prompt-based approaches that do not modify model weights and we compare them with the fine-tuned-based approaches.
    Approach. Following the approach of the study He et al. (2024), we introduced two additional security-oriented prompts, Generic Security and Specific Security, to assess their influence on secure code generation. These were evaluated against the original prompts used in our experiments of RQ1. In the baseline setting, the prompt simply instructed: “Please complete the code”. For the Generic Security setting, we appended the instruction: “Make sure that the generated code is secure, meaning it does not contain any security vulnerabilities”, to enforce the model to generate secure code. The Specific Security setting was further designed to target particular vulnerabilities by including the CWE ID and its description in the prompt template. All code generation settings, including the generation of 30 samples per scenario and a temperature of 0.8, remained consistent with those used in RQ1.
    Prompt Template for the Original One ### Instruction:
    Please complete the code.
    <<Security-Sensitive Scenario>>
    Prompt Template for the Generic Security ### Instruction:
    Please complete the code.
    Make sure that the generated code is secure, meaning it does not contain any security vulnerabilities.
    <<Security-Sensitive Scenario>>
    Prompt Template for the Specific Security ### Instruction:
    Please complete the code.
    Make sure that the generated code is secure with respect to the CWE-<<ID>> vulnerability, meaning it does not contain the following security vulnerability:
    CWE-<<ID>>:<<Description of the CWE>>
    <<Security-Sensitive Scenario>>
    Results. The results are presented in Table 6. The results of LoRA-tuned and IA3-tuned models are the results from RQ1. Overall, the two security-oriented prompts (Generic Security and Specific Security prompts) achieved the highest secure ratio in three out of eight settings. In comparison, fine-tuning-based approaches outperformed in four out of eight settings. On average, the Generic Security prompt attained a secure ratio of 59.9%, while the Specific Security prompt achieved 57.5%. Among the fine-tuned models, the LoRA-tuned models achieved the highest average secure ratio at 60.9%, and IA3-tuned models are 58.5%.
    Table 6: The comparative results of LLMs for secure code generation in prompt-based approaches (Generic and Specific) and PEFT techniques (LoRA-tuned and IA3-tuned).
    Model C C++
    Valid code / Secure code / Valid code / Secure code /
    Total Valid code Total Valid code
    Pretrained 720/780 (92.3%) 403/720 (56.0%) 674/780 (86.4%) 414/674 (61.4%)
    Generic 703/780 (90.1%) 404/703 (57.5%) 670/780 (85.9%) 415/670 (61.9%)
    CodeGen2-7B Specific 711/780 (91.2%) 421/711 (59.2%) 682/780 (87.4%) 401/682 (58.8%)
    LoRA-tuned 711/780 (91.2%) 444/711 (62.4%) 651/780 (83.5%) 393/651 (60.4%)
    IA3-tuned 695/780 (89.1%) 419/695 (60.3%) 694/780 (89.0%) 423/694 (61.0%)
    Pretrained 734/780 (94.1%) 443/734 (60.4%) 703/780 (90.1%) 402/703 (57.2%)
    Generic 734/780 (94.1%) 446/734 (60.8%) 699/780 (89.6%) 403/699 (57.7%)
    CodeLlama-7B Specific 744/780 (95.4%) 453/744 (60.9%) 716/780 (91.8%) 418/716 (58.4%)
    LoRA-tuned 654/780 (83.8%) 397/654 (60.7%) 621/780 (79.6%) 372/621 (59.9%)
    IA3-tuned 734/780 (94.1%) 427/734 (58.2%) 719/780 (92.2%) 447/719 (62.2%)
    Pretrained 753/780 (96.5%) 436/753 (57.9%) 729/780 (93.5%) 445/729 (61.0%)
    Generic 758/780 (97.2%) 452/758 (59.6%) 736/780 (94.4%) 455/736 (61.8%)
    DeepSeekCoder-7B Specific 754/780 (96.7%) 427/754 (56.6%) 739/780 (94.7%) 427/739 (57.8%)
    LoRA-tuned 759/780 (97.3%) 441/759 (58.1%) 716/780 (91.8%) 448/716 (62.6%)
    IA3-tuned 746/780 (95.6%) 426/746 (57.1%) 725/780 (92.9%) 410/725 (56.6%)
    Pretrained 768/780 (98.5%) 455/768 (59.2%) 734/780 (94.1%) 474/734 (64.6%)
    Generic 767/780 (98.3%) 443/767 (57.8%) 754/780 (96.7%) 469/754 (62.2%)
    Magicoder-6.7B Specific 771/780 (98.8%) 396/771 (51.4%) 766/780 (98.2%) 441/766 (57.6%)
    LoRA-tuned 764/780 (97.9%) 452/764 (59.2%) 741/780 (95.0%) 470/741 (63.4%)
    IA3-tuned 763/780 (97.8%) 436/763 (57.1%) 724/780 (92.8%) 434/724 (59.9%)
    Summary of RQ2: By applying the two prompt-based approaches, we found the LoRA-tuned models still perform the best in secure code generation with 60.9% secure ratio on average.
    RQ3: How the granularity of fine-tuning dataset may impact fine-tuning LLMs for secure code generation?
    Motivation. RQ1 investigates only one version of the fine-tuning dataset, i.e., code blocks from source code files that contain vulnerability fixes. RQ1 findings motivate us to explore diverse versions of the same fine-tuning datasets. For instance, coarse-grained datasets, which contain more complete code structures (e.g., full source files), may provide richer semantic context. However, fine-tuned LLMs may learn incomplete semantics (“noise”) from the fine-tuning dataset, which results in a larger number of invalid code compared to the pre-trained LLMs. Conversely, finer-grained datasets, such as those containing only modified lines, may reduce this noise and help the LLMs concentrate on learning vulnerability fix patterns more effectively. Balancing these granularities of datasets may be key to maximizing the benefits of fine-tuning for code-related tasks.
    Approach. We crafted three additional fine-tuning datasets based on granularity from the same set of 4,900 vulnerability-fixing commits. The one used in RQ1 is block-level. In this RQ, we created file-level, function-level, and line-level fine-tuning datasets. The file-level dataset consists of all the files that have modifications in the vulnerability-fixing commits. For the function-level dataset, we first located the modified lines in each commit and then extracted only the functions containing these modified lines. Similarly, we collected the block-level dataset by locating the nearest blocks containing the modified lines. Finally, the line-level dataset consists of all the modified lines in the vulnerability-fixing commits. Note that the four granularities of the fine-tuning sets are all from the same 4,900 commits. We extract all the blocks, functions, or lines relevant to the changed lines for each commit. This RQ aims to explore the impact of different code contexts on fine-tuning performance. Table 7 summarizes the size of the four datasets of different granularity. After constructing the four datasets, we apply these datasets to fine-tune the four models using the two parameter-efficient fine-tuning techniques and full fine-tuning.
    Granularity Total Number of Items Used for Fine-tuning
    file-level 14,622 files
    function-level 32,244 functions
    block-level 79,464 blocks
    line-level 1,240,407 lines
    Results. Table 3 shows the vulnerability evaluation results of the four fine-tuned models in different granularities. In general, we notice that the LLMs fine-tuned by finer-grained (block, function, and line) datasets usually perform better than the file-level (in 21 out of 24 testing settings).
    With LoRA, models fine-tuned at the function-level outperform those fine-tuned at other granularities, achieving the highest secure ratio in 5 out of 8 testing settings. For IA3, the line-level fine-tuned models yield the best results, achieving the highest secure ratio in 6 out of 8 settings. Similarly, in the case of full fine-tuning, the line-level dataset leads to the best performance, attaining the highest secure ratio in 7 out of 8 testing settings. CodeGen2 achieves a 6.4% increase in the secure ratio for C when fine-tuned with LoRA using the block-level dataset, compared to the pre-trained model. CodeLlama also performs well with a 60.7% secure ratio on the block-level dataset for C. For DeepSeek-Coder, the function-level dataset yields the best secure performance with LoRA, achieving secure ratios of 59.5% for C (+1.6%) and 63.3% for C++ (+2.3%) over the pre-trained model. Additionally, the function-level dataset improves the valid ratios compared to the pre-trained baseline. In Magicoder, the file-level dataset delivers the highest secure ratio for C under LoRA, while the function-level and line-level datasets achieve the best results for C++, reaching a secure ratio of 66.0%.
    Similar to the findings of RQ1, full fine-tuning achieves the highest secure ratios compared to LoRA, IA3, and the pre-trained models. However, it also results in a higher number of invalid code snippets. We observed that using the file-level dataset yields the highest valid ratio among all granularities. In contrast, applying finer-grained datasets during full fine-tuning leads to severe catastrophic forgetting, negatively impacting the models’ ability to generate valid code.
    (a) The comparison of the pre-trained models and the fine-tuned models using LoRA with block-level dataset in C language for each CWE.
    (b) The comparison of the pre-trained models and the fine-tuned models using LoRA with function-level dataset in C++ language for each CWE.
    Figure 7: Secure ratio of fine-tuned LLMs for the 11 CWEs.
    Summary of RQ3: Finer-grained vulnerability-fixing code often results in fine-tuned LLMs with better performance in PEFT techniques, as evidenced by a higher secure ratio. In contrast, the file-level dataset makes full fine-tuned models generate more valid code snippets. This suggests that in PEFT techniques, finer-grained datasets may compel LLMs to learn vulnerability fixing patterns more effectively by excluding irrelevant code.
@@
    Approach. Since we focus on C/C++ in this study, we selected tasks of C++ language from HumanEval-X. We call it as HumanEval_CPP. For each task, we generated 10 samples for each pre-trained and fine-tuned version of the four models. Then, we calculate PASS@1, meaning that for one prompt, whether or not an LLM can generate one correct code given one attempt.
    Results. Table 8 presents PASS@1 results of all our fine-tuned LLMs on HumanEval_CPP dataset. CodeGen2 suffers a slight performance degradation, especially when fine-tuned with IA3. CodeLlama shows improvements when fine-tuned by IA3, and comparable results (no performance degradation) when fine-tuned by LoRA. When using file-level in CodeGen2, it achieved 9.09% and 9.33% in LoRA and IA3, respectively, which are close to the performance of the pre-trained model. Additionally, we conducted a Fisher exact test to compare the code generation results between one original LLM and one fine-tuned LLM for all the 164 tasks in the HumanEval dataset. Our results show that for all the LLMs in our study, no significant difference exists between the original and the fine-tuned LLMs. For most of the 164 tasks, the P-value is greater than 0.05, which means that it does not affect the functional correctness of the code generation between the pre-trained model and the fine-tuned model. To give an example, comparing CodeLlama and the fine-tuned CodeLLama using LoRA and file-level fine-tuning dataset, for only two of 164 tasks in HumanEval, we observed a statistically significant difference. DeepSeek-Coder experienced an average 5.0% drop in PASS@1 between the pre-trained model and the model fine-tuned with LoRA. Similarly, Magicoder showed an average decrease of 11.3% in PASS@1. Interestingly, the models fine-tuned using IA3 achieved the highest PASS@1 scores with 25.6% for DeepSeek-Coder and 38.5% for Magicoder, which suggests that the IA3 method aligns well with the task format (i.e., function-based code generation). In contrast, full fine-tuning led to significant performance degradation across all four models on the HumanEval_CPP benchmark.
@@
    Line-level 8.05% Line-level 10.73%
@@
    CodeGen2-7B File-level 9.33% CodeLlama-7B File-level 12.87%
@@
    Full Full
    File-level 2.56% File-level 6.16%
    Function-level 3.78% Function-level 3.84%
    Block-level 3.54% Block-level 2.44%
    Line-level 1.59% Line-level 2.07%
    Pre-trained 24.15% Pre-trained 37.50%
    LoRA LoRA
    File-level 19.27% File-level 25.18%
    Function-level 19.39% Function-level 27.26%
    Block-level 18.78% Block-level 26.52%
    Line-level 19.15% Line-level 25.73%
    IA3 IA3
    DeepSeek-Coder-7B File-level 17.87% Magicoder-6.7B File-level 36.83%
    Function-level 25.61% Function-level 38.54%
    Block-level 21.28% Block-level 36.10%
    Line-level 20.49% Line-level 35.30%
    Full Full
    File-level 14.15% File-level 19.88%
    Function-level 6.83% Function-level 9.82%
    Block-level 4.94% Block-level 6.04%
    Line-level 2.93% Line-level 5.00%